# Tensor products of bimodules and bimodule pairs over monoids

### v2

@@ 3.1. Examples of hypersemigroups and their pairs
Let us first recall some of the famous hyperfields (and their accompanying hyperpairs, cf. Definition 3.4), from [1]. In every case the sum of two distinct nonzero elements is not a singleton.
@@
    The tropical hyperfield consists of the set ${}\mathcal{H}=\mathbb{R}\cup\{-\infty\}$, with $\infty$- ∞ here as the zero element and 00 as the unit, equipped with the addition $a\boxplus b=\{a\}$ if $a>b$, $a\boxplus b=\{b\}$ if $a<b$, and $a\boxplus a=[-\infty,a]$.
@@
  3. (iii)
    The phase hyperfield. Let $S^{1}$ denote the complex unit circle, and take $\mathcal{H}=S^{1}\cup\{0\}$. Nonzero points $a$ and $b$ are antipodes if $a=-b.$ Multiplication is defined as usual (so corresponds on $S^{1}$ to addition of angles). We denote an open arc of less than 180 degrees connecting two distinct points $a,b$ of the unit circle by $({a,b})$. The hypersum is given, for $a,b\neq 0$, by
    $a\boxplus b=\begin{cases}({a,b})\text{ if }a\neq\pm b;\\ \{-a,0,a\}\text{ if }a=-b\enspace,\\ \{a\}\text{ if }a=b\enspace.\end{cases}$
    The hypersystem $\overline{\mathcal{H}}$ contains only the points of $\mathcal{H}$, the elements of the form $a\boxplus b$ with $a,b\in S^{1}$, and subsets $C$ of $\mathcal{H}$, where either $C$ is an open half circle connecting an element $a\in S^{1}$ to $-a$ (which is obtained as the sum $a\boxplus b\boxplus(-a)$, where $b\in C$), or $C=\mathcal{H}$ (which is obtained as $a\boxplus b\boxplus(-a)\boxplus(-b)$).
    $e=\mathbb{1}\boxplus-\mathbb{1}=\{-\mathbb{1},0,\mathbb{1}\},$ so $e\boxplus e=e$. Thus the phase hyperfield is $\circ$-idempotent, but is not $\circ$-distributive since $(a,b)e=(a,b)\cup\{\mathbb{0}\}\cup(-a,-b),$ whereas $(a,b)\boxplus(-a,-b)$ is the whole circle.
@@
    (The weak phase hyperfield [1]; Akian-Gaubert’s modification of the phase hyperfield). The non-zero elements can still be represented by elements of the unit circle $S^{1}$. However, the hyperaddition differs, for $a,b\neq 0$, we now have
    $a\boxplus b=\begin{cases}{[a,b]}&\text{ if }a\neq b\text{ and }a\neq-b,\\ S^{1}\cup\{0\}&\text{ if }a=-b,\\ \{a\}&\text{ if }a=b.\end{cases}$
    where ${[a,b]}$ denotes the closed arc of the angle inferior to $180$ degrees joining $a$ and $b$ (compare with the open arc $(a,b)$ in the phase hyperfield). As the phase hyperfield, the weak phase hyperfield is not distributive.
@@
Let $\mathcal{H}$ be any nonempty set. Stipulate a hyperzero $\mathbb{0}\in\mathcal{H}.$ In each case, its hyperpair $(\mathcal{A},\mathcal{A}_{0})$ will satisfy property N.
@@
      Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\{\mathbb{0},s\}$, $\mathbb{0}\boxplus s=s\boxplus\mathbb{0}=s,$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and $-s=s.$
@@
      (As in [18]) Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}\setminus s$ and $s\boxplus s^{\prime}=\{s,s^{\prime}\}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ This is a hypergroup when $|\mathcal{H}|\geq 3.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup\{\mathcal{H}\}.$ Again $-s=s.$
@@
      Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s^{\prime}=\mathcal{H}\setminus\{s,s^{\prime}\}$ for all nonzero $s,s^{\prime}\in\mathcal{H}.$ This is a hypergroup when $|\mathcal{H}|\geq 4.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup_{s,s^{\prime}\in\mathcal{H}}(H\setminus\{s,s^{\prime}\})\cup\{\mathcal{H}\}.$ (Here $H\setminus\{s\}=(s^{\prime}\boxplus s^{\prime\prime})\boxplus s)$ where $s^{\prime},s^{\prime\prime}$ are distinct from $\{\mathbb{0},s\}.$)
@@
    Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=s$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all $s\neq s^{\prime}$ in $\mathcal{H}.$ This is idempotent, but is not uniquely negated.
@@
      $a_{1}\boxplus a_{2}=\mathcal{H}$. This gives the admissible pair ({},$(\mathcal{H}\cup\{\mathcal{H}\},\{\mathbb{0},\{\mathcal{H}\})$ which satisfies Property N.
@@
    Now suppose $\mathcal{H}$ is ordered, with $\mathbb{0}\in\mathcal{H}$ minimal. Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}$ and $s\boxplus s^{\prime}=s^{\prime}$ for all $\mathbb{0}\neq s<s^{\prime}$ in $\mathcal{H}.$ This is an idempotent hypergroup giving rise to an admissible hyperpair $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and $\mathcal{A}_{0}=\{\mathcal{H}\}$.
In each case, $\mathcal{H}$ becomes a hyper-semiring when $\mathcal{H}$ is a cancellative multiplicative monoid.
@@ 3.2. Residue (quotient) hypermodules and hyperpairs
Suppose that $\mathcal{M}$ is a $C$-bimodule where the ring $C$ is viewed as a multiplicative monoid $\mathcal{T}$ as in Remark 2.3(i), and $G$ is a normal multiplicative subgroup of $\mathcal{T}$, in the sense that $bG=Gb$ for all $a\in\mathcal{M}$. Define the residue hypermodule $\mathcal{H}=\mathcal{M}/G$ over $\mathcal{T}/G$ to have multiplication induced by the cosets, and hyperaddition :×$\boxplus:\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{M})$ by
@@
As defined in Remark 3.2(iv), $(\mathcal{M},\mathcal{M}_{0})$ is a pair, by [1, Proposition 7.21]. For the applications in the hyperfield literature one would take the classical field $\mathcal{M}$ with a multiplicative subgroup $\mathcal{T}$.
@@
As in Remark 3.5, define $b_{1}G\preceq b_{2}G$ when for each $g\in G$ there is $g^{\prime}\in G$ such that $b_{1}g\preceq b_{2}g$. We claim that if $b_{i}G\preceq b_{i}^{\prime}G$ for $i=1,2,$ then $b_{1}G+b_{2}G\preceq b_{1}^{\prime}G+b_{2}^{\prime}G.$
Indeed, if $b=b_{1}g_{1}+b_{2}g_{2}\in b_{1}G+b_{2}G,$ then taking $b_{i}^{\prime}g_{i}^{\prime}\succeq b_{i}g_{i}$ in $b_{i}^{\prime}G,$ we have
$b=b_{1}g_{1}+b_{2}g_{2}\preceq=b_{1}^{\prime}g_{1}^{\prime}+b_{2}^{\prime}g_{2}^{\prime}\in b_{1}^{\prime}G+b_{2}^{\prime}G.$
@@
    The weak phase hyperfield can be obtained by taking the quotient $\mathcal{F}/G$, where =$\mathcal{F}=\mathbb{C}\{\{t^{\mathbb{R}}\}\}$, and $G$ is the group of (generalized) Puiseux series with positive real leading coefficient, where the leading coefficient is the coefficient $f_{\lambda}$ of the series $f=\sum_{\lambda\in\Lambda}f_{\lambda}t^{\lambda}$ such that $\lambda$ is the minimal element of $\{\lambda\in\Lambda:f_{\lambda}\neq 0\}$.
4. Tensor products of bimodules and bimagmas
Remark 4.1.
Let us start with the observation that the classical treatment of a right module and a left module over a ring $C$ only uses the fact that $C$ is a monoid. So we shall talk of tensor products over a multiplicative monoid.
In [12] tensor products were treated from the point of view of semirings, following [13, 26]. There is a delicate issue here, concerning which are the morphisms in our category. One may start with the more straightforward instance of homomorphisms. Then it is rather easy to construct a tensor product satisfying the theory of [9, 21], and their theorems are applicable.
When we bring in weak morphisms on both sides, things become much more complicated, and our approach must become more compatible with [4] and [20].
@@ 4.1. Construction of the tensor product
  1. (i)
    The free magma $(\mathcal{F}(X),*)$ on a set $X$ is the set of formal expressions with parentheses, defined by $X\subseteq\mathcal{F}(X)$ of length 1, and inductively, for $v,w\in\mathcal{F}(X)$ of lengths $m$ and $n$, $(v*w)\in\mathcal{F}(X)$ of length $m+n.$
  2. (ii)
    The free abelian semigroup $(\mathcal{S}(X),+)$ on a set $X$ is the set of formal sums without parentheses, defined by $X\subseteq\mathcal{S}(X)$ of length 1, and inductively, for $v,w\in\mathcal{S}(X)$ of lengths $m$ and $n$, $v+w=w+v\in\mathcal{F}(X)$ of length $m+n.$
We follow the exposition of the classical tensor product, as in [23, Chapter 18]. Throughout this section, $\mathcal{M}_{1}$ is a right $\mathcal{T}$-module and $\mathcal{M}_{2}$ is a left $\mathcal{T}$-module, for convenience.
@@
    Define the $\mathcal{T}$-tensor product semigroup [^10] More generally, in parallel to [20], for $\mathcal{M}_{1}$ a right $\mathcal{T}$-magma and $\mathcal{M}_{1}$ a left $\mathcal{T}$-magma, we can define the $\mathcal{T}$-tensor product magma $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the magma $\mathcal{F}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by all $\bigg{(}\big{(}v_{1}*w_{1},x_{2}\big{)},\big{(}v_{1},x_{2})*(w_{1},x_{2}\big{)}\bigg{)},\quad\bigg{(}\big{(}x_{1},v_{2}*w_{2}\big{)},(x_{1},v_{2})*(x_{1},w_{2})\bigg{)},$ (4.1) $\bigg{(}(x_{1}a,x_{2}),(x_{1},ax_{2})\bigg{)}$ (4.2) $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$. The remainder of this section would go over, mutatis mutandis, but the notation is less familiar. $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the semigroup $\mathcal{S}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by all
@@
    When $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ have negation maps, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big{(}(-)v_{1},v_{2}\big{)},\big{(}v_{1},(-)v_{2})\big{)}$ for all $v_{i}\in\mathcal{F}_{i}.$
@@
    A simple tensor of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is an element of the form $v\otimes w$ for $v\in\mathcal{M}_{1}$ and $w\in\mathcal{M}_{2}$.
@@
    If $a_{1}w_{1}=a_{2}w_{2},$ then $v_{1}a_{1}\otimes w_{1}+v_{2}a_{2}\otimes w_{2}=(v_{1}+v_{2})\otimes a_{2}w_{2}$. This process of “recombining” is straightforward when $\mathcal{T}$ is a group, since then it is applicable for any sum
    $v_{1}\otimes w_{1}+v_{2}\otimes w_{2}=(v_{1}a_{1}^{-1}+v_{2}a_{2}^{-1})\otimes a_{2}w_{2}.$
@@
As in [4], one could define the tensor product $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$ of a right $\mathcal{T}$-monoid $\mathcal{T}_{1}$ and a left $\mathcal{T}$-monoid $\mathcal{T}_{2}$ to be $(\mathcal{T}_{1}\times\mathcal{T}_{2})/\Phi$ where where $\Phi$ is the congruence generated by all $\bigg{(}(x_{1}a,x_{2}),(x_{1},ax_{2})\bigg{)}$, $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$.
4.1.1. Tensor products of prepairs
In order to define the tensor product prepair, we need to define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$.
Definition 4.6.
  1. (i)
    Suppose $(\mathcal{M}_{1},{\mathcal{M}_{1}}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule prepair and $(\mathcal{M}_{2},{\mathcal{M}_{2}}_{0})$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule prepair. Define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}:=({\mathcal{M}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{M}_{2})\cup(\mathcal{M}_{1}\otimes_{\mathcal{T}}{\mathcal{M}_{2}}_{0}).$
  2. (ii)
    Suppose $(\mathcal{M}_{1},{\mathcal{M}_{1}}_{0})$ is a $(\mathcal{T}_{1},\mathcal{T})$-bimodule pair and $(\mathcal{M}_{2},{\mathcal{M}_{2}}_{0})$ is a $(\mathcal{T},\mathcal{T}_{2})$-bimodule pair. Define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}:=({\mathcal{M}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{M}_{2})+(\mathcal{M}_{1}\otimes_{\mathcal{T}}{\mathcal{M}_{2}}_{0}).$

### v4

@@ 3.1. Examples of hypersemigroups and their pairs
Let us first recall some of the celebrated hyperfields (and their accompanying hyperpairs, cf. Definition 3.3), from [1]. In every case we take $\mathcal{T}=\mathcal{H},$ and the hypersum of two distinct nonzero elements is not a singleton.
@@
    The tropical hyperfield consists of the set $\mathcal{H}=\mathbb{R}\cup\{-\infty\}$, with -$\infty$ here as the zero element and 0 as the unit, equipped with the addition $a\boxplus b=\{a\}$ if $a>b$, $a\boxplus b=\{b\}$ if $a<b$, and $a\boxplus a=[-\infty,a]$.
@@
    The hyperfield of signs $L:=\{0,1,-1\}$ has the intuitive multiplication law, and hyperaddition defined by $\boxplus 1=1, 11$$\boxplus-1=-1$, $x\boxplus 0=0\boxplus x=x$ for all $x,$ and $1\boxplus-1=-1\boxplus 1=\{0,1,-1\}.$
  3. (iii)
    The phase hyperfield. Let $S^{1}$ denote the complex unit circle, and take $\mathcal{H}=S^{1}\cup\{0\}$. Nonzero points $a$ and $b$ are antipodes if a=b.a=-b. Multiplication is defined as usual (so corresponds on $S^{1}$ to addition of angles). In this example, we denote an open arc of less than 180 degrees connecting two distinct points $a,b$ of the unit circle by (a,b)({a,b}). The hypersum is given, for $a,b\neq 0$, by
@@
    $\langle\mathcal{T}\rangle$ contains only the points of $\mathcal{H}$, the elements of the form $a\boxplus b$ with $a,b\in S^{1}$, and subsets $C$ of $\mathcal{H}$, where either $C$ is an open half circle connecting an element $a\in S^{1}$ to a-a (which is obtained as the sum $a\boxplus b\boxplus(-a)$, where $b\in C$), or $C=\mathcal{H}$ (which is obtained as $a\boxplus b\boxplus(-a)\boxplus(-b)$).
    $e=\mathbb{1}\boxplus-\mathbb{1}=\{-\mathbb{1},0,\mathbb{1}\},$ so $e\boxplus e=e$. Thus the phase hyperfield is $\mathcal{A}_{0}$-idempotent, but is not $\circ$-distributive since $(a,b)e=(a,b)\cup\{\mathbb{0}\}\cup(-a,-b),$ whereas $(a,b)\boxplus(-a,-b)$ is the whole circle.
@@
    (The weak phase hyperfield [1]; Akian-Gaubert’s modification of the phase hyperfield). The non-zero elements can still be represented by elements of the unit circle $S^{1}$. However, the hyperaddition differs – for $a,b\neq 0$, we now have
[…]
    where [a,b]{[a,b]} denotes the closed arc of less than $180$ degrees joining $a$ and $b$ (compare with the open arc $(a,b)$ in the phase hyperfield). Like the phase hyperfield, the weak phase hyperfield is not distributive.
@@
Let $\mathcal{H}$ be a set with an element stipulated $\mathbb{0}.$ In each of the following cases, $\mathcal{H}$ will be a hypersemigroup whose hyperpair $(\mathcal{A},\mathcal{A}_{0})$ will satisfy property N.
@@
      Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\{\mathbb{0},s\}$, $\mathbb{0}\boxplus s=s\boxplus\mathbb{0}=s,$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and s=s.-s=s.
@@
      (As in [18]) When $|\mathcal{H}|\geq 3,$ define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}\setminus s$ and $s\boxplus s^{\prime}=\{s,s^{\prime}\}$ for all nonzero $s\neq s^{\prime}$ in $\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup\{\mathcal{H}\}.$ Again s=s.-s=s.
@@
      When $|\mathcal{H}|\geq 4,$ define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s^{\prime}=\mathcal{H}\setminus\{s,s^{\prime}\}$ for all nonzero $s,s^{\prime}\in\mathcal{H}.$ Here $\mathcal{A}=\mathcal{H}\cup_{s,s^{\prime}\in\mathcal{H}}\{s,s^{\prime}\}\cup_{s\in\mathcal{H}}(H\setminus\{s\})\cup_{s,s^{\prime}\in\mathcal{H}}(H\setminus\{s,s^{\prime}\})\cup\{\mathcal{H}\}.$ (Here $H\setminus\{s\}=(s^{\prime}\boxplus s^{\prime\prime})\boxplus s)$ where $s^{\prime},s^{\prime\prime}$ are distinct from $\{\mathbb{0},s\}.$)
@@
    Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=s$ and $s\boxplus s^{\prime}=\mathcal{H}$ for all $s\neq s^{\prime}$ in $\mathcal{H}.$ This is idempotent, and is not uniquely negated.
@@
      $a_{1}\boxplus a_{2}=\mathcal{H}$. This gives the admissible pair $(\mathcal{H}\cup\{\mathcal{H}\},\{\mathbb{0},\{\mathcal{H}\}\})$, which satisfies Property N.
@@
    Now take $\mathcal{H}$ ordered, with $\mathbb{0}\in\mathcal{H}$ minimal. Define $\boxplus$ on $\mathcal{H}$ by $s\boxplus s=\mathcal{H}$ and $s\boxplus s^{\prime}=s^{\prime}$ for all $\mathbb{0}\neq s<s^{\prime}$ in $\mathcal{H}.$ Taking $\mathcal{A}=\mathcal{H}\cup\{\mathcal{H}\},$ and $\mathcal{A}_{0}=\{\mathcal{H}\}$, the hyperpair $(\mathcal{A},\mathcal{A}_{0})$ is admissible and $\mathcal{A}_{0}$-bipotent.
In each case above, $\mathcal{H}$ becomes a hyper-semiring when $\mathcal{H}$ is a cancellative multiplicative monoid.
@@ 3.2. Residue (quotient) hypermodules and hyperpairs
Suppose that $\mathcal{M}$ is a $C$-bimodule where the ring $C$ is viewed as a multiplicative monoid $\mathcal{T}$ as in Remark 2.3(i), and $G$ is a normal multiplicative subgroup of $\mathcal{T}$, in the sense that $bG=Gb$ for all $a\in\mathcal{M}$. Define the residue hypermodule $\mathcal{H}=\mathcal{M}/G$ over $\mathcal{T}/G$ to have multiplication induced by the cosets, and hyperaddition $\boxplus:\mathcal{H}\times\mathcal{H}\to\mathcal{P}(\mathcal{M})$ by
@@
As defined in Remark 3.2(iv), $(\mathcal{M},\mathcal{M}_{0})$ is a pair, by [1, Proposition 7.21]. For the applications in the hyperfield literature one would take $\mathcal{M}$ to be a field with a multiplicative subgroup $\mathcal{T}$.
@@
As in Remark 3.4, define $b_{1}G\preceq b_{2}G$ when for each $g\in G$ there is $g^{\prime}\in G$ such that $b_{1}g\preceq b_{2}g$. We claim that if $b_{i}G\preceq b_{i}^{\prime}G$ for $i=1,2,$ then $b_{1}G\boxplus b_{2}G\preceq b_{1}^{\prime}G\boxplus b_{2}^{\prime}G.$
Indeed, if $b=b_{1}g_{1}+b_{2}g_{2}\in b_{1}G\boxplus b_{2}G,$ then taking $b_{i}^{\prime}g_{i}^{\prime}\succeq b_{i}g_{i}$ in $b_{i}^{\prime}G,$ we have
$b=b_{1}g_{1}+b_{2}g_{2}\preceq=b_{1}^{\prime}g_{1}^{\prime}+b_{2}^{\prime}g_{2}^{\prime}\in b_{1}^{\prime}G\boxplus b_{2}^{\prime}G.$
@@
    The weak phase hyperfield can be obtained by taking the quotient $\mathcal{F}/G$, where $\mathcal{F}=\mathbb{C}\{\{t^{\mathbb{R}}\}\}$, and $G$ is the group of (generalized) Puiseux series with positive real leading coefficient, where the leading coefficient is the coefficient $f_{\lambda}$ of the series $f=\sum_{\lambda\in\Lambda}f_{\lambda}t^{\lambda}$ such that $\lambda$ is the minimal element of $\{\lambda\in\Lambda:f_{\lambda}\neq 0\}$.
Lemma 3.12.
If $\mathcal{M}$ is free over $\mathcal{T}$ with a $G$-invariant base $\mathcal{B}$}, then $\bar{\mathcal{M}}:=\mathcal{M}/G$ is free over $\bar{\mathcal{B}}:=\{b_{i}G,b_{i}\in{\mathcal{B}}\}$,
Proof.
We may assume ${\mathcal{B}}=\{b_{i}g_{i}:i\in I\}.$ Clearly $\bar{\mathcal{B}}$ spans $\bar{\mathcal{M}}$. But if $\sum(a_{i}G)b_{i}G=\sum(a^{\prime}_{i}G)b_{i}G,$ then $\sum a_{i}b_{i}^{\prime\prime}g_{i}=\sum a^{\prime}_{i}b_{i}g_{i}^{\prime}g_{i}$ for suitable $g_{i},g_{i}^{\prime},g_{i}^{\prime\prime}\in G,$ implying $a_{i}g_{i}g^{\prime\prime}_{i}=a_{i}^{\prime}g_{i}^{\prime},$ and thus $a_{i}G=a_{i}^{\prime}G.$
4. Tensor products of bimodules
In [12] tensor products were treated from the point of view of semirings, following [13, 26]. There is a delicate issue here, concerning which are the morphisms in our category. One may start with the more straightforward instance of homomorphisms. Then it is rather easy to construct a tensor product satisfying the theory of [9, 21], whose theorems are applicable.
When we bring in weak morphisms on both sides, things become much more complicated, and our approach becomes more compatible with [4] and [20].
@@ 4.1. Construction of the tensor product
The free abelian semigroup $(\mathcal{S}(X),+)$ on a set $X$ is the set of formal sums without parentheses, defined by $X\subseteq\mathcal{S}(X)$ of length 1, and inductively, for $v,w\in\mathcal{S}(X)$ of lengths $m$ and $n$, $v+w=w+v\in\mathcal{S}(X)$ of length $m+n.$
We follow the exposition of the classical tensor product, as in [23, Chapter 18]. Throughout this section, $\mathcal{M}_{1}$ is a right $\mathcal{T}$-module and $\mathcal{M}_{2}$ is a left $\mathcal{T}$-module, for convenience [^10] More generally, for $\mathcal{M}_{1}$ a right $\mathcal{T}$-magma and $\mathcal{M}_{1}$ a left $\mathcal{T}$-magma, we can define the free magma $(\mathcal{F}(X),*)$ on a set $X$ is the set of formal expressions with parentheses, defined by $X\subseteq\mathcal{F}(X)$ of length 1, and inductively, for $v,w\in\mathcal{F}(X)$ of lengths $m$ and $n$, $(v*w)\in\mathcal{F}(X)$ of length $m+n,$ and the $\mathcal{T}$-tensor product magma $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the magma $\mathcal{F}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by $\big(\big(v_{1}*w_{1},x_{2}\big),\big(v_{1},x_{2})*(w_{1},x_{2}\big)\big),\ \big(\big(x_{1},v_{2}*w_{2}\big),(x_{1},v_{2})*(x_{1},w_{2})\big),$ $\big((x_{1}a,x_{2}),(x_{1},ax_{2})\big)$ $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$. The remainder of this subsection would go over, mutatis mutandis, but the notation is more involved..
@@
    Define the $\mathcal{T}$-tensor product semigroup $\mathcal{M}_{1}\otimes_{\mathcal{T}}\mathcal{M}_{2}$ to be the semigroup $\mathcal{S}(\mathcal{M}_{1}\times\mathcal{M}_{2})/\Phi,$ where $\Phi$ is the congruence generated by all
@@
    When $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ have negation maps, we incorporate the negation map into the definition of tensor product, by enlarging $\Phi$ to include also $(\big((-)v_{1},v_{2}\big),\big(v_{1},(-)v_{2})\big)$ for all $v_{i}\in\mathcal{M}_{i}.$
@@
    A simple tensor of $\mathcal{M}_{1}\otimes\mathcal{M}_{2}$ is an element of the form $v_{1}\otimes v_{2}$ for $v_{i}\in\mathcal{M}_{i}.$
@@
    If $a_{1}v_{2}=a_{2}w_{2},$ then $v_{1}a_{1}\otimes v_{2}+w_{1}a_{2}\otimes w_{2}=(v_{1}+w_{1})\otimes a_{2}w_{2}$. This process of “recombining” is straightforward when $\mathcal{T}$ is a group, since then it is applicable for any sum
    $v_{1}\otimes v_{2}+w_{1}\otimes w_{2}=(v_{1}a_{1}^{-1}+w_{1}a_{2}^{-1})\otimes a_{2}w_{2}.$
@@
As in [4], one could define the tensor product $\mathcal{T}_{1}\otimes_{\mathcal{T}}\mathcal{T}_{2}$ of a right $\mathcal{T}$-monoid $\mathcal{T}_{1}$ and a left $\mathcal{T}$-monoid $\mathcal{T}_{2}$ to be $(\mathcal{T}_{1}\times\mathcal{T}_{2})/\Phi$ where where $\Phi$ is the congruence generated by all $\bigg((x_{1}a,x_{2}),(x_{1},ax_{2})\bigg)$, $\forall x_{i}\in\mathcal{\mathcal{M}}_{i},$ $a\in\mathcal{T}$.
4.1.1. Tensor products of bimodule prepairs
In order to define the tensor product bimodule prepair, we need to define $(\mathcal{M}_{1}\otimes\mathcal{M}_{2})_{0}$.